OpenStudy (anonymous):

Let f(x,y)=(2x−y)^5 Find= ∂2f∂x∂y = ∂3f∂x∂y∂x = ∂3f∂x2∂y = Let f(x,y)=(2x−y)^5 Find= ∂2f∂x∂y = ∂3f∂x∂y∂x = ∂3f∂x2∂y = @Mathematics

OpenStudy (amistre64):

wow, they really want you to hunt these dont they

OpenStudy (amistre64):

d/dxdy i believe is forst dx then dy; then to 2 it you go thru it again

OpenStudy (amistre64):

f(x,y)=(2x−y)^5 fx(x,y)=5(2x−y)^4 * 2 = 10(2x-y)^4 fxy(x,y)=40(2x−y)^3 *-1 = -40(2x-y)^3 that should be one round

OpenStudy (anonymous):

Thats correct, thanks!

OpenStudy (anonymous):

I do not understand how to do the next one though, could you list steps for that as well?

OpenStudy (amistre64):

the steps are to go thru it each time from the last results, if that makes sense

OpenStudy (amistre64):

in other words; if the first step is to Fx, then we find Fx(y) from it

OpenStudy (anonymous):

yeah but odnt I use different numbers when taking fx(Y)

OpenStudy (amistre64):

d3f/(dxdydx) this one might mean to find dx, then dy, then dx; 3 times

OpenStudy (amistre64):

in Fsy you "y" it from the Fx that you determined

OpenStudy (amistre64):

in Fxy .. that is

OpenStudy (anonymous):

so I just derive the answer for f(x) ?

OpenStudy (amistre64):

yes, Fxy would mean: derive Fx, then from Fx, derive a "y" to get Fxy

OpenStudy (amistre64):

partials mean we consider any variable we are not respecting, as a constant

OpenStudy (amistre64):

Fx means we consider all ys to be constants

OpenStudy (anonymous):

-240(2x-y)^2 is the answer, you have helped a bunch...thanks!

OpenStudy (amistre64):

F(x,y)=(2x−y)^5 \[F_x = 5(2x-y)^4*(2x-y)^{'x}\] \[F_x = 5(2x-y)^4*2\] whew!! good :)

OpenStudy (amistre64):

\[F_x = 10(2x-y)^4\] \[F_{xy} = 10*4(2x-y)^3\ (2x-y)^{'y}\] \[F_{xy} = 10*4(2x-y)^3\ (-1)\] \[F_{xy} = -40(2x-y)^3\] \[F_{xyx} = -40*3(2x-y)^2\ (2x-y)^{'x}\] \[F_{xyx} = -40*3(2x-y)^2\ (2)\] \[F_{xyx} = -240(2x-y)^2 \]